jeremiah23forver:

Roger pushes a box on a 30° incline. If he applies a force of 60 newtons parallel to the incline and displaces the box 10 meters along the incline, how much work will he do on the box?

Shadow:

\[W = F \times D \times \cos \theta\]

Shadow:

You legit just have to fill in the numbers.